OpenStudy (anonymous):

What is 3 3/8 · 4 1/3 ?

OpenStudy (ranga):

Convert each mixed number to a fraction: 3 3/8 = ((3*8)+3)/8 = 27/8 4 1/3 = ((4*3)+1)/3 = 13/3 Multiply the two fractions: 27/8 * 13/3 = (27*13)/(8*3) The 3 at the bottom will go 9 times into the 27 at the top. (9*13)/8 = 117/8 = 14 5/8